OpenStudy (anonymous):

Let a = 6 What is the value of 4 +2a?

OpenStudy (dan815):

4+2*6=?

OpenStudy (dan815):

4+2*a a=6 4+2*6=?

OpenStudy (yadu123):

dan can you help me in probability

OpenStudy (anonymous):

16